OpenStudy (anonymous):

Show all steps used to solve the problem. One point for steps and one point for solution. Be sure that you check your solution for correctness. 17-x/4=-10

OpenStudy (lgbasallote):

one good step here is to bring the x on the other side so it's positive 17 = x/4 - 10 now isolate x 17+10 = x/4 27 = x/4 what do you think comes next?

OpenStudy (anonymous):

times 27 with 4 =108 ?

OpenStudy (lgbasallote):

yeahh